# SHuBERT: Self-Supervised Sign Language Representation Learning via Multi-Stream Cluster Prediction

### v1

@@ 4.4 Phonological Feature Recognition
We also conduct experiments on phonological feature recognition, that is the classification of linguistic features of signs, for two of the ISLR datasets. We report the Recall at 1 (prediction accuracy) for 16 commonly used phonological features (from ASL-LEX 2.0 [46]) in Tab. 7. The training setups and hyperparameters are identical to those of the full fine-tuning method in Sec. 4.3, except that we now train 16 classification heads simultaneously. We also find that removing weight decay gives a slight performance boost.
@@
To the best of our knowledge, no previous work has reported phonological feature recognition accuracies on the ASL Citizen dataset. Similarly, though the Sem-Lex authors [26] report phonological feature prediction accuracies, they are not comparable to ours, which are computed on the entire test set available to the public. Thus, we hope that our results in Tab. 7 can serve as a benchmark for future work.
4.5 Ablations
Masking Strategies
Masking strategy BLEU-1 BLEU BLEURT
Channel masking 14.5 2.6 29.9
Time masking 15.1 2.3 31.2
Random masking 15.4 2.2 31.4
Tab. 8 compares translation performance when using each of the three masking strategies (Fig. 4). We observe different signals from different metrics. While the BLEU scores suggest channel masking to be the best, random masking produces better BLEURT scores. Contradicting signals between different evaluation metrics for sign language translation has also been observed in prior work [63]. We chose to priotize BLEURT, as it generally has better alignment with human judgements of translation quality [15], and therefore use random masking in all of our other experiments.
Data Scaling Behavior of SHuBERT
Table 9: Impact of pre-training data size on SHuBERT’s performance.
Hours of pre-train data BLEU-1 BLEU BLEURT
984 15.4 2.2 31.4
98 12.7 0.7 29.1
In addition to the full pre-training dataset, we also train SHuBERT from scratch on a randomly selected 10% of the full data. In Tab. 9 we see that there is a noticeable drop in performance in BLEU and BLEURT, suggesting that data size is important. With multiple larger datasets now available—BOBSL[1], JWSign[18] and YouTube-SL-25[54] contain approximately 1500, 2500, and 3200 hours of data respectively—we expect that expanding the SHuBERT pre-training data may further improve performance. In addition, these larger datasets also include more language diversity, which may also improve performance and/or applicability to additional languages.
Isolating SHuBERT’s Impact on Performance
Layer of SHuBERT BLEU-1 BLEU BLEURT
None 15.3 2.5 31.6
Last Layer 21.4 4.7 35.0
Weighted Sum 29.3 7.1 39.5
To understand SHuBERT’s impact on translation performance, we conduct three experiments, shown in Tab. 10. In our baseline experiment (“None”), we directly feed the projected 4-channel features (face, left hand, right hand, body pose) to the ByT5 translation model, bypassing SHuBERT entirely, resulting in fairly poor performance. When we instead pass these features through a frozen pre-trained SHuBERT (400k steps, random masking) and use its final layer’s output (“Last Layer”), we see significant improvement. Finally, computing a learned weighted combination of all SHuBERT layers (“Weighted Sum”) further improves performance. These results demonstrate that each component of SHuBERT contribute to translation performance.
Frozen vs. Fine-Tuned SHuBERT
Fine-tune SHuBERT BLEU-1 BLEU BLEURT
$\times$ 21.4 4.7 35.0
$\checkmark$ 30.0 7.5 39.9
We also investigate the impact of fine-tuning during translation training, as shown in Tab. 11. We compared two scenarios: using a frozen SHuBERT (400k steps, random masking) and only fine-tuning ByT5, versus fine-tuning both SHuBERT (from the same base model) and ByT5 together. Both scenarios use features from SHuBERT’s last layer. Fine-tuning SHuBERT leads to substantial improvements compared to keeping it frozen, when using the final layer. However, referring back to Tab. 10, the relatively small difference between the fine-tuned performance and that of the frozen and layer-weighted SHuBERT is noteworthy. This observation is promising for low-resource sign languages, where we may have plentiful unlabeled video data for pre-training, but very limited parallel data for translation training.
5 Conclusion
SHuBERT, our proposed self-supervised approach for learning sign language video representations, yields a transformer encoder that maps from multiple feature streams (face and hand appearance and upper body pose) to a stream of per-frame contextual representations. A single base SHuBERT model, when adapted to a range of sign language processing benchmarks including both translation and isolated sign recognition, achieves strong performance on all of them and almost always improves over the state of the art. SHuBERT is trained on public data and will be released upon publication. Based on its strong performance on the tasks studied here, we expect that SHuBERT can serve as a base model for a broad range of sign language processing tasks. In addition to the directions for future work mentioned in the relevant experiment sections above, SHuBERT has the potential for further improvement using techniques like iterative re-clustering and re-training, such as those used for similar speech models [20].
Limitations
Our work has several limitations. First, although the results are competitive with or outperform prior work, the absolute performance is still quite poor. Neither our model nor others can replace human interpreters for broad-domain sign language translation. Second, our training data volume is significantly smaller than that of typical self-supervised speech and text models. We cannot say with certainty how our model would scale up to much larger datasets. Finally, the current scope of our work is limited to American Sign Language. Although sign languages use the same channels and share many elements, we do not know how well our model would generalize to other languages. Future work could address these limitations by expanding the training dataset and training on data from other sign languages.
Acknowledgment
Supplementary Material
6 ASL Phonological Feature Classification Details
@@
American Sign Language (ASL) can be described through a set of phonological features, similarly to the description of spoken languages via features. These features capture the essential components of sign formation, including hand configuration, movement patterns, and spatial relationships. Tab. 12 presents the number of classes for each of the phonological features we use in our phonological feature prediction analysis for Sem-lex and ASL-Citizen, and below we list the values of each feature. This commonly used feature set is from ASL-LEX 2.0 [46]).
@@ 7 Cluster Samples
Figure 7: Sample upper body clusters. Each row represents a distinct cluster and 10 random examples from it. Cluster 5 seems to correspond to configurations of the upper body pose where the right hand is at shoulder level and the left hand is down. Cluster 43 seems to be a configuration where the two hands are raised and close to each other near the chest, and the signer is facing slightly to the right. This might correspond to signs being performed with both hands being involved/active. Cluster 36 appears similar, but with the hands slightly farther apart. Cluster 19 is generally similar to cluster 43, except that the signer tends to be facing slightly to the left. Finally in cluster 168, the right hand is usually above the shoulder and close to the face/head.
8 Translation examples
We provide example translations produced by our model fiven inputs from three ASL datasets: How2Sign (instructional content) Tab. 13, FLEURS-ASL (zero-shot setting) Tab. 14, and OpenASL (general domain with native signers and varying background) Tab. 15.
@@ 8 Translation examples
Uthus et al. [55] It’s really great for women’s self defense.
Rust et al. [44] This is a really great point for women’s self defense.
@@
Uthus et al. [55] In this clip we’re going to show you how to cut a piece of clay.
Rust et al. [44] In this clip I’m going to show you how to clip the cable, the cable.
@@
Uthus et al. [55] In this clip we’re going to talk about how to feed a set of baiting lizards for a lava field oil.
Rust et al. [44] In this clip we’re going to talk about how to feed the trail for draining clean for laborer oil.
@@
Uthus et al. [55] Their hopping and dancing is now, they’re going to need their squat and squat and they’re going to be able to move independently.
Rust et al. [44] So that she’s going to get her hips up as far as she can, and now she’s going to lift her head up as far as possible.
@@
Uthus et al. [55] But if you have to set up a new campfire, there are two ways to do a low impact fire, one is a cone fire, which we have to do in the tent earlier, and the other one is to set up a campfire in a fire pan.
Rust et al. [44] But if you have to set up a new campfire, this is one way to do it in a low impact. One is a monk fire. One is a campfire. The other one is to set a campfire in a campfire. That’s just a post like the top of the post.
@@
Uthus et al. [55] Alright, let’s get started.
Rust et al. [44] It’s an important part of the process.
@@
Table 14: Qualitative translation examples from the FLEURS-ASL dataset, comparing SHuBERT-based translation (zero-shot) to a previous approach [53].
@@
Tanzer [53] In the 1980s, she worked in theaters like taxesi, cheesy, and tracy.
@@
Tanzer [53] There is a new technique to detect brains and vision.
@@
Tanzer [53] The law requires all states to agree on a standard and that it is a legal requirement.
@@
Table 15: Qualitative translation examples from the OpenASL dataset, comparing SHuBERT-based translations to a previous model [50]
. (1) Reference thank you Shi et al. [50] thank you Ours thank you (2) Reference come on Shi et al. [50] come on Ours maybe (3) Reference now i’ve come this far and it ’s a different team Shi et al. [50] how do you feel about it Ours it feels like a crash in the team (4) Reference i was there from the beginning to the end and time went by fast Shi et al. [50] the students were thrilled by this Ours i just wanted to leave because it went ahead and started (5) Reference i’m here at nad’s 50th wow Shi et al. [50] the nad has been ¡unk¿ for many years Ours well the nad is shocked to have 50 years of dhh (6) Reference i entered the yap 2018 competition and won Shi et al. [50] the competition was started with ideas Ours i enrolled in that competition in 2018 and then i won (7) Reference you can check out their kickstarter in the link below Shi et al. [50] you can watch the conversation at lake county Ours you can check out their kickstarter link below (8) Reference that is one thing i found interesting and wanted to share with you today Shi et al. [50] i also am the president of the jr. nad conference here Ours that’s one interesting thing she wanted to share with you (9) Reference those are the different types of bills Shi et al. [50] schools have switched to teaching students Ours i looked at several different types of interpreting services (10) Reference dry january has picked up in popularity since it began in 2012 Shi et al. [50] krispy kreme is bringing back its original playstation in 2016 Ours the qury dry january started in 2012 (11) Reference we will be happy to respond give you support and listen to your concerns Shi et al. [50] please review and submit your time passion and support this important issue Ours the nad is willing to respond and support your concerns (12) Reference there were videos posted on the internet that showed a person walking on the grass completely engulfed in flames Shi et al. [50] a video shows the officer walking up to his shoulder and before he was shot Ours videos posted on social media of him walking on a grass walking completely with fire (13) Reference and people would become carpenters laborers mechanics plowers and farmers Shi et al. [50] the next year 1880 the nad was established in the first operation 30 of the house in 2015 Ours and he was forced to wear a wearing a labover meganic and a financial warper (14) Reference for nad youth programs related information please contact us via facebook at the nad youth programs or email us through Shi et al. [50] you can contact us through our website where you can check our facebook page online at ¡unk¿ Ours if you want to contact the nad youth program you can contact us through our facebook page at the nad youth program or youth program through our website (15) Reference last week suspects gregory mcmichael and his son travis were arrested and charged with felony murder and aggravated assault Shi et al. [50] last week a black man named ¡unk¿ ¡unk¿ was arrested and charged with felony murder and aggravated assault Ours last week two suspects gregory mcmichael and his son travis were arrested and charged with felony murder and wounded by another gravated assaulter

### v3

@@ Appendix B Phonological Feature Recognition
Table 9: Phonological feature recognition accuracy on two datasets, Sem-Lex and ASL Citizen.
@@
We also conduct experiments on phonological feature recognition, that is the classification of linguistic features of signs, for two of the ISLR datasets. We report the Recall at 1 (prediction accuracy) for 16 commonly used phonological features (from ASL-LEX 2.0 Sevcikova Sehyr et al. (2021)) in Tab. 9. The training setups and hyperparameters are identical to those of the full fine-tuning method in Sec. 4.3, except that we now train 16 classification heads simultaneously. We also find that removing weight decay gives a slight performance boost.
@@
To the best of our knowledge, no previous work has reported phonological feature recognition accuracies on the ASL Citizen dataset. Similarly, though the Sem-Lex authors Kezar et al. (2023) report phonological feature prediction accuracies, they are not comparable to ours, which are computed on the entire test set available to the public. Thus, we hope that our results in Tab. 9 can serve as a benchmark for future work.
Appendix C ASL Phonological Feature Classification Details
@@
American Sign Language (ASL) can be described through a set of phonological features, similarly to the description of spoken languages via features. These features capture the essential components of sign formation, including hand configuration, movement patterns, and spatial relationships. Tab. 10 presents the number of classes for each of the phonological features we use in our phonological feature prediction analysis for Sem-lex and ASL-Citizen, and below we list the values of each feature. This commonly used feature set is from ASL-LEX 2.0 Sevcikova Sehyr et al. (2021)).
@@ Appendix D Cluster Samples
Figure 8: Sample upper body clusters. Each row represents a distinct cluster and 10 random examples from it. Cluster 5 seems to correspond to configurations of the upper body pose where the right hand is at shoulder level and the left hand is down. Cluster 43 seems to be a configuration where the two hands are raised and close to each other near the chest, and the signer is facing slightly to the right. This might correspond to signs being performed with both hands involved/active. Cluster 36 appears similar, but with the hands slightly farther apart. Cluster 19 is generally similar to cluster 43, except that the signer tends to be facing slightly to the left. Finally in cluster 168, the right hand is usually above the shoulder and close to the face/head.
Appendix E Translation examples
We provide example translations produced by our model given inputs from three ASL datasets: How2Sign (instructional content, Tab. 11), FLEURS-ASL (zero-shot setting, Tab. 12), and OpenASL (general domain with native signers and varying background, Tab. 13).
@@ Appendix E Translation examples
Uthus et al. (2023) It’s really great for women’s self defense.
Rust et al. (2024) This is a really great point for women’s self defense.
@@
Uthus et al. (2023) In this clip we’re going to show you how to cut a piece of clay.
Rust et al. (2024) In this clip I’m going to show you how to clip the cable, the cable.
@@
Uthus et al. (2023) In this clip we’re going to talk about how to feed a set of baiting lizards for a lava field oil.
Rust et al. (2024) In this clip we’re going to talk about how to feed the trail for draining clean for laborer oil.
@@
Uthus et al. (2023) Their hopping and dancing is now, they’re going to need their squat and squat and they’re going to be able to move independently.
Rust et al. (2024) So that she’s going to get her hips up as far as she can, and now she’s going to lift her head up as far as possible.
@@
Uthus et al. (2023) But if you have to set up a new campfire, there are two ways to do a low impact fire, one is a cone fire, which we have to do in the tent earlier, and the other one is to set up a campfire in a fire pan.
Rust et al. (2024) But if you have to set up a new campfire, this is one way to do it in a low impact. One is a monk fire. One is a campfire. The other one is to set a campfire in a campfire. That’s just a post like the top of the post.
@@
Uthus et al. (2023) Alright, let’s get started.
Rust et al. (2024) It’s an important part of the process.
@@
Tanzer (2024b) In the 1980s, she worked in theaters like taxesi, cheesy, and tracy.
@@
Tanzer (2024b) There is a new technique to detect brains and vision.
@@
Tanzer (2024b) The law requires all states to agree on a standard and that it is a legal requirement.
@@
Table 12: Qualitative translation examples from the FLEURS-ASL dataset, comparing SHuBERT-based translation (zero-shot) to a previous approach Tanzer (2024b).
(1) Reference thank you
Shi et al. (2022) thank you
Ours thank you
(2) Reference come on
Shi et al. (2022) come on
Ours maybe
(3) Reference now i’ve come this far and it ’s a different team
Shi et al. (2022) how do you feel about it
Ours it feels like a crash in the team
(4) Reference i was there from the beginning to the end and time went by fast
Shi et al. (2022) the students were thrilled by this
Ours i just wanted to leave because it went ahead and started
(5) Reference i’m here at nad’s 50th wow
Shi et al. (2022) the nad has been <unk> for many years
Ours well the nad is shocked to have 50 years of dhh
(6) Reference i entered the yap 2018 competition and won
Shi et al. (2022) the competition was started with ideas
Ours i enrolled in that competition in 2018 and then i won
(7) Reference you can check out their kickstarter in the link below
Shi et al. (2022) you can watch the conversation at lake county
Ours you can check out their kickstarter link below
(8) Reference that is one thing i found interesting and wanted to share with you today
Shi et al. (2022) i also am the president of the jr. nad conference here
Ours that’s one interesting thing she wanted to share with you
(9) Reference those are the different types of bills
Shi et al. (2022) schools have switched to teaching students
Ours i looked at several different types of interpreting services
(10) Reference dry january has picked up in popularity since it began in 2012
Shi et al. (2022) krispy kreme is bringing back its original playstation in 2016
Ours the qury dry january started in 2012
(11) Reference we will be happy to respond give you support and listen to your concerns
Shi et al. (2022) please review and submit your time passion and support this important issue
Ours the nad is willing to respond and support your concerns
(12) Reference there were videos posted on the internet that showed a person walking on the grass completely engulfed in flames
Shi et al. (2022) a video shows the officer walking up to his shoulder and before he was shot
Ours videos posted on social media of him walking on a grass walking completely with fire
(13) Reference and people would become carpenters laborers mechanics plowers and farmers
Shi et al. (2022) the next year 1880 the nad was established in the first operation 30 of the house in 2015
Ours and he was forced to wear a wearing a labover meganic and a financial warper
(14) Reference for nad youth programs related information please contact us via facebook at the nad youth programs or email us through
Shi et al. (2022) you can contact us through our website where you can check our facebook page online at <unk>
Ours if you want to contact the nad youth program you can contact us through our facebook page at the nad youth program or youth program through our website
(15) Reference last week suspects gregory mcmichael and his son travis were arrested and charged with felony murder and aggravated assault
Shi et al. (2022) last week a black man named <unk> <unk> was arrested and charged with felony murder and aggravated assault
Ours last week two suspects gregory mcmichael and his son travis were arrested and charged with felony murder and wounded by another gravated assaulter
Table 13: Qualitative translation examples from the OpenASL dataset, comparing SHuBERT-based translations to a previous model Shi et al. (2022)